# Med-R2: Crafting Trustworthy LLM Physicians via Retrieval and Reasoning of Evidence-Based Medicine

### v4

@@ Appendix A Evidence-Based Medicine (EBM)
Evidence-Based Medicine (EBM) is defined as the conscientious, explicit, and judicious application of the best current evidence in making decisions regarding the care of individual patients. This practice entails the integration of individual clinical expertise with the most reliable external clinical evidence derived from systematic research [40]. The EBM process typically includes five stages, question formulating, evidence searching, evidence appraising, evidence applying and effect assessing, which aims to make the best possible health care decision through iterative improvements.
To utilize “best evidence”, researchers assess the quality of trials by determining the grading system based on the likelihood that the methods used and the results obtained are less prone to bias and more reliable. The hierarchy of evidence guides the clinical decision-making, since not all evidence is created equal, as described in Figure 6. The evidence hierarchy establishes the priority of references, particularly when conflicting facts are present within the retrieved evidence.
Figure 6: Illustration for Hierarchy of evidence. The base of the pyramid represents the lowest quality and highest risk of bias in research design, while the apex signifies the highest quality and lowest risk of bias.
The construction and rationale behind the hierarchical structure of evidence grading is outlined from the highest to the lowest levels:
  • Systematic Reviews/Meta-Analyses (SR/MA) represent the highest tier of evidence. These assessments evaluate the consistency and risk of bias across all research findings within the medical domain, demonstrating the overall effect of interventions or exposures.
@@
  • Individual Case Reports are of the second-lowest evidence level, essentially uncontrolled cohort studies lacking a comparison group. The absence of a control group affects the correlation between study variables and interventions/exposures/risk factors and outcomes.
  • Expert Opinion is considered the lowest level of evidence due to its high susceptibility to bias. Compared to other levels, experts are more likely to choose evidence that confirms their preconceived hypotheses or beliefs, potentially leading to conflicts of interest and a focus on a specific domain while overlooking broader contexts, thereby introducing bias into their perspectives.
@@ Appendix B Method Details
B.1 Question Formulation Details
B.1.1 Details of Query Classification
@@
Evidence-Based Medicine (EBM) Categories  Non-professional queries may fail to clearly articulate the current medical symptoms, thereby hindering the effective retrieval of the necessary evidence. Here we classify medical queries into several types according to the categories of Evidence-Based Medicine (EBM) questions, including diagnosis, therapy, prognosis, etiology, prevention, cost, etc., and then conduct professional medical query reformulations based on their EBM categorization to emphasize specialized retrieval. Instructions for targeted augmentation of clinical queries are outlined in Table 4, where we employ the Qwen2.5-72B-Instruct [^7] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct model fine-tuned on our 100 manually annotated samples as our question reformulator.
General Natural Language Question Types  It serves as a crucial reference for filtering and reranking the evidence documents retrieved in response to a query, since the emphasis of the expected answer varies with different question types. We categorize queries into 12 natural language classes, as illustrated in Table 3, establishing a mapping between the types of questions and the types of retrieved evidence documents to form a question-answer typology (depicted in Figure 8).
Table 3: Explanations and descriptions for each category within the general question classification.
@@
We have categorized the question of each instance within the datasets utilized for our experiments. For detailed statistical values of the data samples, please refer to Table 7. We employed the complete set of instances (“# Total Instance” in Table 7) from each dataset for the classification process. The 3D bar charts of the classification results are shown in Figure 3 and Figure 7.
(a) Query category of MedQA-USMLE.
(b) Query category of MedQA-MCMLE.
(c) Query category of PubMedQA.
(d) Query category of MMLU-Med.
Figure 7: Query categories of different medical datasets. We employ a 3D bar chart to represent the number of instances in each category within the two distinct classification systems of the current dataset. A logarithmic scale (base 10) on the z-axis, ranging from 1 to $z\,(z=100,140,4000)$, is utilized to represent the wide range of values.
Table 4: Prompts for query reformulation of each category within the Evidence-Based Medicine (EBM) categories.
Category Instructions for Query Reformulation
B.1.2 Details of Query Reformulator
Following the instructions outlined in Table 4, we have manually annotated 120 samples (20 samples for each EBM category) to fine-tune Qwen2.5-72B-Instruct, and employed the trained model to professionally reformulate the queries within the medical datasets to enhance the subsequent retrieval of relevant evidence pertaining to the current query from the knowledge corpus.
B.2 Knowledge Corpus Details
B.2.1 Data Sources of Medical Knowledge Base
We organized the medical knowledge corpus from four data sources: academic papers, entries, books, and guidelines. List below are the data sources we built upon:
  • Academic Papers  Academic literature provide valuable insights derived from the latest scientific investigations, which offers a robust theoretical foundation for guiding clinical practice and informing public health decisions. We obtained the publicly available literature from PubMed Central (PMC) [^8] https://pmc.ncbi.nlm.nih.gov/, and exracted valid contents following the processing pipeline of  Hakala et al. [14].
@@ B.2.1 Data Sources of Medical Knowledge Base
  • Books  Medical textbooks are vital resources for medical knowledge retrieval, which can be consulted when faced with complex cases or when seeking updated knowledge about specific conditions. We gathered books from NCBI Bookshelf [^10] https://www.ncbi.nlm.nih.gov/books/, and finally collected 10k documents following the handling process of  Hakala et al. [14].
  • Guidelines  Clinical practice guidelines serve as essential tools in Evidence-Based Medicine (EBM), aiding healthcare providers and patients in making informed decisions regarding diagnosis and treatment. We utilized the guideline data when training MEDITRON series from  Chen et al. [7], and gained approximately 10k documents.
The statistical information on volume and token number across different knowledge resources can be found in Table 1. Based on our analysis, 95.8% of the documents in the entire medical knowledge corpus have a length of less than 10,000 tokens. Consequently, we set the threshold for content segmentation at 10,000 tokens.
B.2.2 Details of the Retrieval Strategy
Motivation for Choosing Hybrid Dense and Sparse Retrieval  The hybrid dense and sparse retrieval leverages the strengths of both dense and sparse representations to enhance retrieval performance, which is particularly beneficial in the medical domain [33]. Dense retrieval (e.g., DPR [26], Sentence-BERT [37]) captures the contextual and deep semantic meanings, while sparse retrieval (e.g., BM25 [38], TF-IDF [36]) excels at accurately matching specific keywords. Such strategy achieves a harmonious integration of semantic comprehension and precise keyword alignment. Prior to the fine-grained reranking process, a substantial number of documents are retrieved and await detailed reordering. Consequently, the goal during the evidence retrieval phase is to identify documents that can assist in addressing the current medical question as comprehensively as possible, which renders hybrid dense and sparse retrieval the most suitable option. In Table 5 we present the strengths and weaknesses of different retrieval strategies:
Table 5: Comparison among different retrieval strategies.
Strategy Semantic Understanding Keyword Precision Recall vs. Precision
Dense Retrieval Strong Weak High recall rate
Sparse Retrieval Weak Strong High precision
Hybrid Dense and Sparse Retrieval Strong Strong Balances both
B.3 Evidence Reranking Details
B.3.1 Components of Fine-Grained Reranker
@@
  • Usefulness $f_{u}(x)$: 1) Regarding the lightweight proxy models we employed for scoring usefulness $f_{u}$ of the retrieved documents, for the LLaMA series, we used the Sheared-LLaMA-1.3B [48] as a lightweight reference model; for the Qwen series, we utilized Qwen-2.5-1.5B [51]. 2) Moreover, to avoid potential label leakage when computing $f_{u}$ in Equation 3, we have implemented strategies to ensure that the entire framework does not have access to ground truth answers (labels) during the deployment phase, where we employed frontier models’ responses to serve as the “reference answers” for consultation throughout the framework’s operation. Specifically, we utilized the training subsets from MedQA-USMLE, MedQA-MCMLE [21], and MedMCQA [35] to fine-tune the Qwen2.5-72B-Instruct [^11] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct model, resulting in model $M_{F}$. We replaced the labels with responses generated by model $M_{F}$ for the current medical question.
@@
Table 6: Explanations and descriptions for each category within the general document classification.
@@
Figure 8: Query Document Projection.
Output: Scores based on document category $f_{g}(S^{c})$
Define $q^{c}$: category of the query used for retrieving
Define $\vec{p}$: category probability distribution of document $d$
1:  for each document $d_{i}$ in $\mathcal{S}^{c}$ do
2:     /* Step 1: Expected Category Mapping */
4:     /* Step 2: Document Category Probability Inference */
5:     Provide category probability distribution of $d_{i}$ referring to $M_{P}$:             $\vec{p_{i}}=\{p(d_{i}|c_{j})\}_{j=1}^{|C|}\leftarrow M_{P}(d_{i})$
6:     /* Step 3: Statistics Aggregation */
8:  end for
9:  Return $f_{g}(S^{c})=\left\{f_{g}(d_{i})\right\}_{i=1}^{|S^{c}|}$
B.3.2 Explanation for Equation Derivation
@@
B.4 Evidence Assessment Details
Here, we provide the details of the iterative multi-document retrieval process, as illustrated in Algorithm 2. There are several points that require clarification.
@@ B.4 Evidence Assessment Details
  • Accuracy Calculation  To avoid potential label leakage during deployment, we also employed the responses of frontier models to serve as the “reference answers” for consultation throughout the operation (line 9 in Algorithm 2). This implementation is the same as the strategy when computing the usefulness score $f_{u}(x)$ described in Section B.3.
Parameter: Coarsely reranked documents $\mathcal{S}^{c}$, finely reranked documents $\mathcal{S}^{f}$
Define $E(d)$: embedding of document $d$
1:  $\mathcal{S}_{\mathcal{Q}-top}^{f},\mathcal{Q}^{CoT}\leftarrow[\,],\,[\,]$
2:  for each query $q_{i}$ in $\mathcal{Q}$ do
3:     for $t=1,2,\ldots,T$ do
5:        /* Step 1: CoT Generation */
6:        Select top $k$ documents from $\mathcal{S}_{i}^{f^{(t)}}$ referring to $w$: $\mathcal{S}_{i-top}^{f^{(t)}}\leftarrow\text{SelectTopK}(\mathcal{S}_{i}^{f^{(t)}},k,w)$
7:        Generate chain-of-thought sequence based on $q_{i}$ and $\mathcal{S}_{i-top}^{f^{(t)}}$:             $q^{CoT^{(t)}}_{i}=\leftarrow M_{CoT}(q_{i},\mathcal{S}_{i-top}^{f^{(t)}})$
8:        /* Step 2: Evidence Assessment */
9:        Compute answer accuracy: $Acc^{(t)}_{i}\leftarrow\text{CalculateAccuracy}(q^{CoT^{(t)}}_{i})$
10:        Compute semantic stability of documents retrieved in consecutive iterations:       $\vec{\mu}^{(t)}_{i}\leftarrow||\frac{1}{|\mathcal{S}_{i}^{c^{(t)}}|}\sum_{d_{i}\in\mathcal{S}_{i}^{c^{(t)}}}E(d_{i})-\frac{1}{|\mathcal{S}_{i}^{c^{(t-1)}}|}\sum_{d_{i}\in\mathcal{S}_{i}^{c^{(t-1)}}}E(d_{i})||$
11:        /* Step 3: Termination Condition Evaluation */
12:        if $Acc^{(t)}_{i}/\vec{\mu}^{(t)}_{i}>\delta$ then
13:           $\mathcal{S}_{\mathcal{Q}-top}^{f}.append(\mathcal{S}_{i-top}^{f^{(t)}})$, $\mathcal{Q}^{CoT}.append(q^{CoT^{(t)}}_{i})$
14:           Break
15:        end if
16:     end for
17:     $\mathcal{S}_{\mathcal{Q}-top}^{f}.append(\mathcal{S}_{i-top}^{f^{(T)}})$, $\mathcal{Q}^{CoT}.append(q^{CoT^{(T)}}_{i})$
18:  end for
19:  Return Selected evidence documents $\mathcal{S}_{\mathcal{Q}-top}^{f}$ and CoT sequences $\mathcal{Q}^{CoT}$
@@ Appendix C Experiments Details
C.1 Details of Medical Datasets
To evaluate the performance of our proposed Med-R2 in medical scenarios, we chose five knowledge-intensive medical question-answering tasks for comparison with other approaches, the statistical description of which is provides in Table 7.
  • MedQA-USMLE and MedQA-MCMLE [21]: The MedQA dataset is derived from the professional medical board exams, covering languages of English, simplified Chinese, and traditional Chinese. In this study, we employ the subsets collected from US and Mainland China. The MedQA-USMLE dataset includes text materials from 18 English medical textbooks, while the MedQA-MCMLE dataset is constructed from 33 simplified Chinese medical textbooks.
  • MedMCQA [35]: The MedMCQA dataset is a large-scale multi-choice question answering data corpus that covers 2,400 healthcare topics and 21 medical subjects. Topics of MedMCQA span from medicine (endocrinology, infection, haematology, respiratory, etc.), surgery (general surgery, endocrinology, breast, and vascular surgery, etc.) to radiology & biochemistry, which are sourced from both real-world scenarios and simulated examinations.
  • PubMedQA [22]: The PubMedQA dataset is a novel biomedical question answering dataset originated from PubMed abstracts. To ensure equity, we exclude the origin-given contexts during the retrieval phase. Here we choose the 1,000 expert-annotated instances for downstream task evaluation. PubMedQA is included only in our test set due to the limited number of instances.
  • MMLU-Med [16]: MMLU-Med is a set of six medical tasks (anatomy, clinical knowledge, professional medicine, human genetics, college medicine, college biology) derived from the MMLU dataset. MMLU-Med is also included only in our test set due to its limited quanity.
For the MedQA-USMLE and MedQA-MCMLE datasets, the original data is divided into three parts: train, dev, and test. We utilize the training part directly for model fine-tuning. The dev and test subsets are merged for evaluation, from which we selected 10 (or 11) instances as Chain-of-Thought (CoT) demonstration examples. Since the test portion of MedMCQA does not provide ground truth answers, we use the development set for evaluation. We employ MedQA-USMLE, MedQA-MCMLE, and MedMCQA datasets for within-dataset fine-tuning. For instance, we train with the training subset of MedQA-USMLE and subsequently evaluate with its corresponding test partitions. For cross-dataset fine-tuning setting, we utilize PubMedQA and MMLU-Med. Specifically, we train on MedMCQA and assess performance on PubMedQA and MMLU-Med to test the model’s generalizability across disparate data sources.
Table 7: Details of the experimental datasets. We utilize MedQA-USMLE, MedQA-MCMLE and MedMCQA for within-dataset fine-tuning, while employing PubMedQA and MMLU-Med for cross-dataset fine-tuning.
@@ C.1 Details of Medical Datasets
MedQA-USMLE [21] 10178 2535 10 12723
MedQA-MCMLE [21] 27400 6840 11 34251
MedMCQA [35] 182822 4170 13 187005
PubMedQA [22] - 990 10 1000
MMLU-Med [16] - 1080 9 1089
C.2 Details of Baselines
@@
  • Frontier Models: Frontier models are considered to represent the pinnacle of performance across various dimensions of LLMs, serving as the strongest baselines. Here, we have selected several state-of-the-art LLMs to establish the maximum potential of model performance on several medical benchmarks, including GPT-4o [17], Claude3.5-Sonnet [3], and DeepSeek-V3 [30].
@@
  • Open-Sourced Medical Models: These models stand for the domain-specific models that were specifically trained on medical data. We have selected PMC-LLaMA-7B, PMC-LLaMA-13B [47], MEDITRON-7B and MEDITRON-70B [7] to represent the open-sourced medical models for comparison to assess Med-R2’s relative advantage compared to specialized medical AI systems.
@@
  • Direct Response: In this case, we employ the base model to directly respond to medical queries without aid of additional training on any dataset or augmentation from external knowledge bases.
@@
  • Vanilla RAG [28]: It represents the most traditional and fundamental strategy for utilizing external knowledge bases to assist models in answering questions. In this study, we employ the same medical knowledge base as Med-R2 , but directly utilizing the raw text retrieved by FAISS and combining it with the original query to test the model’s performance on medical tasks.
@@
  • Within-Dataset Fine-Tuning: During the within-dataset training, where the training and test sets originate from the same distribution, models fine-tuned on this data naturally exhibit strong performance gains due to the homogeneity of the datasets. Consequently, both Med-R2 and fine-tuning methods perform comparably well in this setting.
@@
  • Cross-Dataset Fine-Tuning: Cross-dataset training poses a greater challenge by testing the model’s ability to generalize across different distributions. This scenario demands robust transfer learning capabilities, which are significantly bolstered by the integration of rich external medical knowledge bases. Med-R2 excels in this context by effectively retrieving and extracting relevant information from these resources, thereby enhancing its performance on diverse datasets.
@@
  • LLM-AMT [46]: LLM-AMT is a RAG system specifically designed for clinical question answering. It incorporates common RAG components such as a query augmentor, textbook retriever, knowledge refiner, and an LLM reader. LLM-AMT leverages a collection of medical textbooks as an external indexable medical knowledge base. However, in this comparison, we aim to assess the RAG’s capability to retrieve, filter, and apply evidence text from the same external knowledge base. To control variables, we employ the LLM-AMT’s process but replace the externally indexable medical knowledge base with our own constructed medical retrieval corpus.
C.3 Hyper-Parameters Setting

### v5

@@ Appendix A Evidence-Based Medicine (EBM)
Evidence-Based Medicine (EBM) is defined as the conscientious, explicit, and judicious application of the best current evidence in making decisions regarding the care of individual patients. This practice entails the integration of individual clinical expertise with the most reliable external clinical evidence derived from systematic research (Sackett et al., 1996). The EBM process typically includes five stages, question formulating, evidence searching, evidence appraising, evidence applying and effect assessing, which aims to make the best possible health care decision through iterative improvements. To utilize “best evidence”, researchers assess the quality of trials by determining the grading system based on the likelihood that the methods used and the results obtained are less prone to bias and more reliable. The hierarchy of evidence guides the clinical decision-making, since not all evidence is created equal, as described in Figure 4. The evidence hierarchy establishes the priority of references, particularly when conflicting facts are present within the retrieved evidence. The construction and rationale behind the hierarchical structure of evidence grading is outlined from the highest to the lowest levels:
  • Systematic Reviews/Meta-Analyses (SR/MA) represent the highest tier of evidence. These assessments evaluate the consistency and risk of bias across all research findings within the medical domain, demonstrating the overall effect of exposures.
@@
  • Individual Case Reports are of the second-lowest evidence level, essentially uncontrolled cohort studies lacking a comparison group. The absence of a control group affects the correlation between study variables, interventions factors and outcomes.
  • Expert Opinion is considered the lowest level of evidence due to its high susceptibility to bias. Compared to other levels, experts tend to choose evidence that confirms their preconceived hypotheses, potentially leading to conflicts of interest and a focus on a specific domain while overlooking broader contexts, thereby introducing bias into their perspectives.
@@ Appendix B Method Details
B.1. Question Formulation Details
B.1.1. Details of Query Classification
@@
Evidence-Based Medicine (EBM) Categories  Nonprofessional queries may fail to clearly articulate the current medical symptoms, thereby hindering the effective retrieval of the necessary evidence. We classify medical queries into several types according to the categories of Evidence-Based Medicine (EBM) questions, including diagnosis, therapy, prognosis, etiology, prevention, cost, etc., and then conduct professional medical query reformulations based on their EBM categorization to emphasize specialized retrieval. Instructions for targeted augmentation of clinical queries are outlined in Table 4, where we employ Qwen2.5-72B-Instruct [^7] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct fine-tuned on our 100 manually annotated samples as our question reformulator.
General Natural Language Question Types  It serves as a crucial reference for filtering and reranking the documents retrieved in response to a query, since the emphasis of the expected answer varies with different question types. We categorize queries into 12 natural language classes, as illustrated in Table 5, establishing a mapping between the types of questions and that of evidence to form a question-answer typology (depicted in Figure 5).
(a) Query category of MedQA-USMLE.
(b) Query category of MedQA-MCMLE.
(c) Query category of PubMedQA.
(d) Query category of MMLU-Med.
Figure 8. Query categories of different medical datasets. We employ a 3D bar chart to represent the number of instances in each category within the two distinct classification systems of the current dataset. A logarithmic scale (base 10) on the z-axis, ranging from 1 to z(z=100,140,4000)z\,(z=100,140,4000), is utilized to represent the wide range of values.
B.1.2. Details of Query Reformulator
Following the instructions outlined in Table 4, we have manually annotated 120 samples (20 samples for each EBM category) to fine-tune the Qwen2.5-72B-Instruct model, and employed the trained model to professionally reformulate the queries within the medical datasets to enhance the subsequent retrieval of relevant evidence pertaining to the current query from the knowledge corpus.
B.2. Knowledge Corpus Details
We organized the medical knowledge corpus from four data sources: academic papers, entries, books, and guidelines, as listed below:
Table 4. Prompts for query reformulation of each category within the Evidence-Based Medicine (EBM) categories.
Category Instructions for Query Reformulation
Table 5. Explanations and descriptions for each category within the general question classification.
@@ B.2. Knowledge Corpus Details
  • Academic Papers  Academic literature provide valuable insights derived from the latest scientific investigations, offering a robust theoretical foundation for guiding clinical practice and informing public health decisions. We obtained the publicly available literature from PubMed Central (PMC) [^8] https://pmc.ncbi.nlm.nih.gov/, and exracted valid contents following the processing pipeline of  Hakala et al. (2016).
@@
  • Books  Medical textbooks are vital resources for medical knowledge retrieval, which can be consulted when faced with complex cases or when seeking updated knowledge about specific conditions. We gathered books from NCBI Bookshelf [^10] https://www.ncbi.nlm.nih.gov/books/, and collected 10k documents following the process of  Hakala et al. (2016).
  • Guidelines  Clinical practice guidelines serve as essential tools in Evidence-Based Medicine (EBM), aiding healthcare providers in making informed decisions for diagnosis and treatment. We utilized the guideline data when training MEDITRON series from  Chen et al. (2023), and gained approximately 10k documents.
B.3. Evidence Reranking Details
B.3.1. Components of Fine-Grained Reranker
@@
  • Usefulness $f_{u}(x)$: 1) Regarding the lightweight proxy models we employed for scoring usefulness $f_{u}$ of the retrieved documents, for the LLaMA series, we used the Sheared-LLaMA-1.3B (Xia et al., 2023) as a lightweight reference model; for the Qwen series, we utilized Qwen-2.5-1.5B (Yang et al., 2024). 2) Moreover, to avoid potential label leakage when computing $f_{u}$ in Equation 3, we have implemented strategies to ensure that the entire framework does not have access to ground truth answers (labels) during the deployment phase, where we employed frontier models’ responses to serve as the “reference answers” for consultation throughout the framework’s operation. Specifically, we utilized the training subsets from MedQA-USMLE, MedQA-MCMLE (Jin et al., 2020), and MedMCQA (Pal et al., 2022) to fine-tune the Qwen2.5-72B-Instruct [^11] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct model, resulting in model $M_{F}$. We replaced the labels with responses generated by model $M_{F}$ for the current medical question.
@@
Table 6. Explanations and descriptions for each category within the general document classification.
@@
B.3.2. Explanation for Equation Derivation
@@
B.4. Evidence Assessment Details
Here, we provide the details of the iterative multi-document retrieval process, as illustrated in Algorithm 2.
@@ Appendix C Experiments Details
C.1. Details of Medical Datasets
For MedQA-USMLE and MedQA-MCMLE, the original data is divided into three parts: train, dev, and test. We utilize the training part directly for model fine-tuning. The dev and test subsets are merged for evaluation, from which we selected 10 (or 11) instances as Chain-of-Thought (CoT) demonstration examples. Since the test portion of MedMCQA does not provide ground truth answers, we use the development set for evaluation. We employ MedQA-USMLE, MedQA-MCMLE, and MedMCQA datasets for within-dataset fine-tuning. For instance, we train with the training subset of MedQA-USMLE and subsequently evaluate with its corresponding test partitions. We utilize the remaining datasets for cross-dataset fine-tuning setting to test the model’s generalizability on disparate data sources.
Table 7. Overall statistics of datasets. We utilize MedQA-USMLE, MedQA-MCMLE and MedMCQA for within-dataset fine-tuning, while others for cross-dataset fine-tuning.
@@ C.1. Details of Medical Datasets
MedQA-USMLE (Jin et al., 2020) 10178 2535 10 12723
MedQA-MCMLE (Jin et al., 2020) 27400 6840 11 34251
MedMCQA (Pal et al., 2022) 182822 4170 13 187005
PubMedQA (Jin et al., 2019) - 990 10 1000
MMLU-Med (Hendrycks et al., 2021) - 1080 9 1089
NEJMQA (Katz et al., 2024) - 655 10 655
MedXpertQA (Zuo et al., 2025) - 2455 10 2455
RareArena (THUMedInfo, 2025) - 72,661 10 72,661
C.2. Details of Baselines
@@
  • Frontier Models: Frontier models are considered to represent the pinnacle of performance across various dimensions of LLMs, serving as the strongest baselines. Here, we have selected several state-of-the-art LLMs to establish the maximum potential of model performance on several medical benchmarks, including GPT-4o (Hurst et al., 2024), Claude3.5-Sonnet (Anthropic, 2024), and DeepSeek-V3 (Liu et al., 2024).
@@
  • Open-Sourced Medical Models: These models stand for the domain-specific models that were specifically trained on medical data. We have selected PMC-LLaMA-7B, PMC-LLaMA-13B (Wu et al., 2024), MEDITRON-7B and MEDITRON-70B (Chen et al., 2023) to represent the open-sourced medical models for comparison to assess Med-R2’s relative advantage compared to specialized medical AI systems.
@@
  • Direct Response: We employ the base model to directly respond to medical queries without aid of additional training on any dataset or augmentation from external knowledge bases.
@@
  • Vanilla RAG (Lewis et al., 2020): It represents the most traditional and fundamental strategy for utilizing external knowledge bases to assist models in answering questions. In this study, we employ the same medical knowledge base as Med-R2 , but directly utilizing the raw text retrieved by FAISS and combining it with the original query to test the model’s performance on medical tasks.
@@
  • Within-Dataset Fine-Tuning: The training and test sets originate from the same distribution, and models exhibit strong performance gains due to the homogeneity of the datasets.
@@
  • Cross-Dataset Fine-Tuning: It poses a greater challenge by testing the model’s ability to generalize across different distributions, demanding robust transfer learning capabilities, which are bolstered by the integration of rich external medical knowledge.
@@
  • LLM-AMT (Wang et al., 2024): LLM-AMT is a RAG system specifically designed for clinical question answering. It incorporates common RAG components such as a query augmentor, textbook retriever, knowledge refiner, and an LLM reader. LLM-AMT leverages a collection of medical textbooks as an external indexable medical knowledge base. However, in this comparison, we aim to assess the RAG’s capability to retrieve, filter, and apply evidence from the same external knowledge base. To control variables, we employ the LLM-AMT’s process but replace the medical knowledge base with our own constructed medical retrieval corpus.
C.3. Hyper-Parameters Setting